An oxide of chromium is made up of 5.20 g chromium and 5.60 g oxygen. What is the empirical formula of the oxide? (Note: An oxide of nitrogen contains 63.1% oxygen and has a molar mass of 76.0 g/mol.)

1
Expert's answer
2021-12-02T00:22:02-0500

5.2/76= 0.068mol

Molar mass of O= 15.99

5.6/15.99= 0.35mol

63.1/100× 0.068= 0.04291

0.068-0.04291= 0.02509

0.39291/0.02509= 16

= CrO16